OpenStudy (anonymous):

Some birds and squirrels were on a branch of a tree. There were 15 animals on the branch with a total of 50 animal feet. How many birds and how many squirrels were on the branch? Set up the problem. Explain why you set it up that way. Provide the solution.

OpenStudy (anonymous):

b=2 s=4 xs+yb=50. x+y=15. birds(b)have 2 legs. squirrels(s)have 4 legs. x birds plus y squirrels=50 legs there are 15 animals total(x+y) 15xs=60. 10xs=40. 40+10=50, s=10 yb=10, y=5. 10 squirrels, 5 birds. Simple enough.:)

OpenStudy (anonymous):

thanks :)